# Dual-Modeling Decouple Distillation for Unsupervised Anomaly Detection

### v1

@@ 3.4.1. Foreground-aware Anomaly Synthesis
Textures Carpet 91.6/-/87.9 -/98.8/95.8 98.9/98.9/97.0 100/99.2/97.7 -/96.1/- - 95.9/99.5/98.3 99.96/99.26/98.11
Grid 81.0/-/95.2 -/99.0/96.6 100/99.3/97.6 100/99.3/97.7 -/99.1/- - 99.9/99.4/97.7 100/99.46/97.57
Leather 88.2/-/94.5 -/99.3/98.0 100/99.4/99.1 100/99.4/99.2 -/99.7/- - 63.6/99.7/98.7 100/99.76/99.50
Tile 99.1/-/94.6 -/97.4/92.1 99.3/95.6/90.6 99.7/96.6/92.4 -/98.0/- - 99.7/96.8/90.3 100/99.53/97.43
Wood 97.7/-/91.1 -/97.2/93.6 99.2/95.3/90.9 99.3/95.8/93.3 -/97.7/- - 99.6/95.2/93.2 99.74/98.02/96.37
Average 91.5/-/92.7 -/98.3/95.2 99.5/97.7/95.0 99.8/98.1/96.1 -/98.1/- - 91.7/98.1/95.6 99.94/99.21/97.80
Objects Bottle 99.0/-/93.1 -/98.8/95.1 100/98.7/96.6 100/98.8/97.0 -/99.2/- - 99.9/98.7/95.6 100/98.99/97.07
Cable 86.2/-/81.8 -/95.5/87.7 95.0/97.4/91.0 99.2/98.4/93.9 -/97.3/- - 98.4/95.7/87.5 98.43/97.97/92.10
Capsule 86.1/-/96.8 -/98.3/92.2 96.3/98.7/95.8 99.0/98.8/96.4 -/99.1/- - 99.8/97.8/89.9 99.60/99.01/97.06
Hazelnut 93.1/-/96.5 -/98.5/94.3 99.9/98.9/95.5 100/99.2/96.3 -/99.6/- - 99.5/98.6/96.1 99.93/99.28/96.92
Metal_nut 82.0/-/94.2 -/97.6/94.5 100/97.3/92.3 100/98.1/93.0 -/98.6/- - 86.9/97.8/93.2 99.90/98.40/95.25
Pill 87.9/-/96.1 -/97.8/96.5 96.6/98.2/96.4 98.4/98.3/97.0 -/98.7/- - 99.7/98.6/94.9 98.15/99.33/97.77
Screw 54.9/-/94.2 -/98.3/93.0 97.0/99.6/98.2 98.9/99.7/98.6 -/98.5/- - 85.3/96.9/85.6 96.19/99.32/97.32
Toothbrush 95.3/-/93.3 -/98.9/92.2 99.5/99.1/94.5 100/99.1/94.2 -/99.3/- - 94.0/99.1/91.8 100/99.41/95.81
Transistor 81.8/-/66.6 -/82.5/69.5 96.7/92.5/78.0 98.5/94.3/81.8 -/89.1/- - 100/99.2/97.6 99.17/95.63/85.87
Zipper 91.9/-/95.1 -98.5/95.2 98.5/98.2/95.4 98.6/98.8/96.3 -/99.1/- - 99.6/97.9/93.0 99.87/99.41/97.77
Average 85.8/-/90.8 -/96.5/90.9 98.0/97.9/93.4 99.2/98.4/94.5 -/97.9/- - 96.3/98.0/92.5 98.94/98.49/95.10
Total Average 87.7/-/91.4 95.5/97.0/92.1 98.5/97.8/93.9 99.4/98.3/95.0 98.6/97.9/- 99.6/98.2/94.5 94.8/98.1/93.6 99.40/98.85/96.13
@@ 4.3. Main Results
bracket black 90.36/98.14/92.35 89.10/98.05/92.59 - 95.74/98.57/97.18
bracket brown 92.76/97.08/95.11 95.02/97.15/94.58 - 98.94/97.31/96.11
bracket white 87.44/99.32/97.65 90.11/99.43/97.25 - 98.33/99.66/99.29
connector 100/99.45/96.90 100/99.29/95.87 - 100/99.68/98.28
metal plate 100/99.09/96.09 100/99.08/96.15 - 100/99.08/96.51
tubes 96.06/99.12/97.23 94.52/99.13/97.39 - 95.56/99.44/98.58

### v2

@@ 3.4.1. Foreground-aware Anomaly Synthesis
Textures Carpet 91.6/-/87.9 -/98.8/95.8 98.9/98.9/97.0 100/99.2/97.7 98.9/96.1/93.6 99.6/99.1/97.5 95.9/99.5/98.3 99.96/99.26/98.11
Grid 81.0/-/95.2 -/99.0/96.6 100/99.3/97.6 100/99.3/97.7 99.7/99.1/96.4 100/99.2/96.9 99.9/99.4/97.7 100/99.46/97.57
Leather 88.2/-/94.5 -/99.3/98.0 100/99.4/99.1 100/99.4/99.2 100/99.7/99.0 100/99.5/99.2 63.6/99.7/98.7 100/99.76/99.50
Tile 99.1/-/94.6 -/97.4/92.1 99.3/95.6/90.6 99.7/96.6/92.4 100/98.0/95.5 100/95.7/91.1 99.7/96.8/90.3 100/99.53/97.43
Wood 97.7/-/91.1 -/97.2/93.6 99.2/95.3/90.9 99.3/95.8/93.3 97.1/97.7/96.1 99.5/95.3/91.2 99.6/95.2/93.2 99.74/98.02/96.37
Average 91.5/-/92.7 -/98.3/95.2 99.5/97.7/95.0 99.8/98.1/96.1 99.1/98.1/96.1 99.8/97.8/95.2 91.7/98.1/95.6 99.94/99.21/97.80
Objects Bottle 99.0/-/93.1 -/98.8/95.1 100/98.7/96.6 100/98.8/97.0 100/99.2/96.6 100/98.8/97.1 99.9/98.7/95.6 100/98.99/97.07
Cable 86.2/-/81.8 -/95.5/87.7 95.0/97.4/91.0 99.2/98.4/93.9 97.8/97.3/86.4 99.2/98.3/93.4 98.4/95.7/87.5 98.43/97.97/92.10
Capsule 86.1/-/96.8 -/98.3/92.2 96.3/98.7/95.8 99.0/98.8/96.4 97.0/99.1/94.2 98.8/98.8/96.2 99.8/97.8/89.9 99.60/99.01/97.06
Hazelnut 93.1/-/96.5 -/98.5/94.3 99.9/98.9/95.5 100/99.2/96.3 99.9/99.6/97.6 100/99.1/95.7 99.5/98.6/96.1 99.93/99.28/96.92
Metal_nut 82.0/-/94.2 -/97.6/94.5 100/97.3/92.3 100/98.1/93.0 99.5/98.6/95.0 100/97.2/90.8 86.9/97.8/93.2 99.90/98.40/95.25
Pill 87.9/-/96.1 -/97.8/96.5 96.6/98.2/96.4 98.4/98.3/97.0 97.2/98.7/95.3 98.3/98.3/96.6 99.7/98.6/94.9 98.15/99.33/97.77
Screw 54.9/-/94.2 -/98.3/93.0 97.0/99.6/98.2 98.9/99.7/98.6 93.6/98.5/92.5 99.1/99.6/98.2 85.3/96.9/85.6 96.19/99.32/97.32
Toothbrush 95.3/-/93.3 -/98.9/92.2 99.5/99.1/94.5 100/99.1/94.2 99.9/99.3/94.0 100/98.9/92.2 94.0/99.1/91.8 100/99.41/95.81
Transistor 81.8/-/66.6 -/82.5/69.5 96.7/92.5/78.0 98.5/94.3/81.8 98.5/89.1/85.7 100/96.4/85.3 100/99.2/97.6 99.17/95.63/85.87
Zipper 91.9/-/95.1 -98.5/95.2 98.5/98.2/95.4 98.6/98.8/96.3 100/99.1/97.4 99.3/98.5/95.9 99.6/97.9/93.0 99.87/99.41/97.77
Average 85.8/-/90.8 -/96.5/90.9 98.0/97.9/93.4 99.2/98.4/94.5 98.3/97.9/93.5 99.5/98.4/94.1 96.3/98.0/92.5 98.94/98.49/95.10
Total Average 87.7/-/91.4 95.5/97.0/92.1 98.5/97.8/93.9 99.4/98.3/95.0 98.6/97.9/94.4 99.6/98.2/94.5 94.8/98.1/93.6 99.40/98.85/96.13
@@ 4.3. Main Results
bracket black 90.36/98.14/92.35 89.10/98.05/92.59 91.2/97.8/94.5 95.74/98.57/97.18
bracket brown 92.76/97.08/95.11 95.02/97.15/94.58 95.2/96.3/95.2 98.94/97.31/96.11
bracket white 87.44/99.32/97.65 90.11/99.43/97.25 92.7/98.8/97.3 98.33/99.66/99.29
connector 100/99.45/96.90 100/99.29/95.87 100/99.4/96.4 100/99.68/98.28
metal plate 100/99.09/96.09 100/99.08/96.15 100/99.1/95.2 100/99.08/96.51
tubes 96.06/99.12/97.23 94.52/99.13/97.39 96.2/99.2/97.3 95.56/99.44/98.58